# The Cohen–Lenstra moments over function fields via the stable homology of non-splitting Hurwitz spaces

### v2

@@ A.3. A scanning map
Next, we produce well behaved topological models for the source and target of § 4.5. The following definition will be the key ingredient in constructing the models:
@@
Let $M$ be a set with a right action of $\operatorname{Hur}^{X}$ and $N$ be a set with a left action of $\operatorname{Hur}^{X}$. Consider the space $B[M,\operatorname{Hur}^{X},N]$ consisting of points which are of the form
@@
    A finite collection of pairwise disjoint open balls $U_{1},\ldots,U_{n}$ in contained in the interior of $[d_{1},d_{2}]\times[0,1]$.
@@
    A homotopy class of paths $\phi$ from the center of the balls $U_{i}$, viewed as an element of $\mathrm{confbig}_{n}$, to the base point.
@@
    None of the points in $x$ lie in $[d_{1},d_{2}]\times[0,1]-\cup_{1}^{n}U_{i}$, and there is a unique point from $x$ in each $U_{i}$.
@@
    Cutting the element of $\operatorname{hurbig}^{X}$ twice to restrict it to the interval $[d_{1},d_{2}]$ yields a point $y^{\prime}\in\operatorname{hurbig}^{X}$ (see § A.2). Then, using the homotopy class of $\phi$, the corresponding tuple of elements in $X$ associated to $y^{\prime}$ is $\alpha_{1},\ldots,\alpha_{n}$.
@@
    Define $y_{1}\in\operatorname{hurbig}^{X}$ to be the element of $\operatorname{hur}^{X}$ obtained by cutting and restricting to the interval $[0,d_{1}]$, and let $y_{2}$ be the element similarly obtained by using the interval $[d_{2},1]$. We then require that $ay_{1}=a^{\prime}$ and $y_{2}b=b^{\prime}$.
@@
[…]
We next construct a homotopy equivalence showing that the topological space $B[M,\operatorname{Hur}^{X},N]$ models the two-sided bar construction $M\otimes_{\operatorname{hur}^{X}}N$, which we define next. Our proof is an adaptation of the proof of [BDPW23, Theorem 5.2.3]. Indeed, the case $M=N=X=*$ corresponds to the result there.
@@
We produce a map $\sigma:M\otimes_{\operatorname{hur}^{X}}N\to B[M,\operatorname{Hur}^{X},N]$, as follows: Given data as in (A.2) for $H=\operatorname{hur}^{X}$, and using notation from § A.2, let $t_{i}:=t(x_{i})\in\mathbb{R}_{>0}$ denote the time associated to $x_{i}\in\operatorname{hur}^{X}$. We multiply $x_{1},\ldots,x_{n}$ to give an element $x\in\operatorname{hur}^{X}$ with $t:=t(x)=\sum_{i=1}^{n}t_{i}$. First, extend the interval on which $x$ is defined from $[0,t]\times[0,1]$ to $[-\frac{1}{2},t+\frac{1}{2}]\times[0,1]$. We then define $t^{\prime}:=\sum_{i=1}^{n}y_{i}\cdot(\sum_{j=1}^{i}t_{j})$. At this point, it will be useful to recall the cutting construction from § A.2. Choose $\epsilon>0$ small enough so that there are no points of the configuration $x$ in $(t^{\prime}-\frac{1}{2},t^{\prime}-\frac{1}{2}+\epsilon]\times[0,1]$ and $[t^{\prime}+\frac{1}{2}-\epsilon,t^{\prime}+\frac{1}{2})\times[0,1]$. We can then cut $x$ into three pieces, $w,x^{\prime\prime},z\in\operatorname{hurbig}^{X}$ coming from the intervals $[-\frac{1}{2},t^{\prime}-\frac{1}{2}+\epsilon]$, $[t^{\prime}-\frac{1}{2}+\epsilon,t^{\prime}+\frac{1}{2}-\epsilon],[t^{\prime}+\frac{1}{2}-\epsilon,t+\frac{1}{2}]$. Extend $x^{\prime\prime}$ to be of length $1$ by extending the length of the interval on both ends by $\epsilon$, and let $x^{\prime}$ denote the resulting element of $\operatorname{hurbig}^{X}$. Note that $x^{\prime}$ doesn’t depend on $\epsilon$, and neither does the class of $w,z$ in $\pi_{0}\operatorname{Hur}^{X}$.
@@
Having shown $\sigma$ is continuous, we next claim that $\sigma$ is surjective on path components. Every point in $B[M,\operatorname{Hur}^{X},N]$ has a path to a point where the configuration on $[0,1]\times[0,1]$ is empty, by linearly pushing the configuration towards $0\times[0,1]$ on the boundary, and multiplying once points collide with the boundary. A point with an empty configuration is determined by $(a,b)\in M\times N$. This point is the image of the point $(a,b,(1),())\in M\otimes_{\operatorname{hur}^{X}}N$, showing the claimed surjectivity on path components. We view these points as above with empty configuration as giving a copy of $M\times N\subset M\otimes_{\operatorname{hur}^{X}}N$.
@@
We turn to proving $(i)$, and assume that we have a map $f$ as given in the statement of $(i)$. We set up notation to define the map $g$ described in $(i)$. Using compactness, we can find a finite open cover $D^{v}=\cup_{\lambda\in\Lambda}U_{\lambda}$ such that there is an $r>0$ and for each $\lambda\in\Lambda$, there is a $t_{\lambda}\in(0,1)$ such that for all $u\in U_{\lambda}$, the configuration associated to $f(u)$ has no points in $(t_{\lambda}-r,t_{\lambda}+r)\times[0,1]$. Choose a partition of unity $w_{\lambda}$ subordinate to the cover $U_{\lambda}$. Let $u$ be a point in $U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$, and not in $U_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{1},\ldots,\lambda_{n}\}$. Assume that $t_{\lambda_{0}}<t_{\lambda_{1}}<\cdots<t_{\lambda_{n}}$ (which can always be achieved by decreasing $r$ and making all $t_{\lambda}$ distinct). Then we consider the $n$-tuple of elements $x_{1},\ldots,x_{n}\in\operatorname{hur}^{X}$, where $x_{i}$ is obtained from the element of $\operatorname{hur}$ associated to $f(u)$ by cutting along $t_{\lambda_{i-1}}\times[0,1]$ and $t_{\lambda_{i}}\times[0,1]$, taking the configuration with points lying in $[t_{\lambda_{i-1}},t_{\lambda_{i}}]\times[0,1]$, and rescaling the intervals by $\frac{1}{2r}$ so that the time function $t$ applied to this new configuration is $(t_{\lambda_{i}}-t_{\lambda_{i-1}})\frac{1}{2r}$. Define $u^{\prime}$ and $u^{\prime\prime}\in\operatorname{hurbig}^{X}$ to be the elements obtained respectively by cutting to restrict to the intervals $[0,t_{\lambda_{0}}]$ and $[t_{\lambda_{n}},1]$ respectively. We define the map $g:D^{v}\to M\otimes_{\operatorname{hur}^{X}}N$ by sending $u$ to the tuple
@@
Choose a basic open set $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$ containing $f(u)$ and a small open neighborhood $V$ containing $u$ with $V\subset f^{-1}(\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime}))\cap U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$. We may shrink $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$, to assume $d_{1}\leq t_{\lambda_{0}},d_{2}\geq t_{\lambda_{n}}$, where, here, shrinking this neighborhood may involve increasing the number of $U_{\lambda}$ so that there is one such open containing each point of the configuration $x$, appearing in the data for $f(u)$, in $(d_{1},d_{2})\times[0,1]$. By further shrinking $V$, we may assume the functions $w_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{0},\ldots,\lambda_{n}\}$ vanish on $V$. Therefore, using identifications $(1)-(3)$ of the bar construction, for the purposes of proving continuity, we may assume our open cover consists solely of $U_{\lambda_{0}},\ldots,U_{\lambda_{n}}$, i.e, we can ignore the other $U_{\lambda}$. At this point, condition (3) of § A.3 enables us to further shrink $V$ so that ($\star$) holds for all $g(v)$ with $v\in V$. We have therefore found an open set in $D^{v}$ containing $u$ and contained in $g^{-1}(W)$ for $W$ an open set containing $g(u)$. This proves that $g$ is continuous.
@@
To prove $(ii)$, given a map of pairs $g:(S^{v},*)\to(M\otimes_{\operatorname{hur}^{X}}N,M\times N)$ and a nullhomotopy of $\sigma\circ g$, the same construction used above for $(i)$ lifts the nullhomotopy of $\sigma\circ g$ to a pointed map $h:D^{v+1}\to M\otimes_{\operatorname{hur}^{X}}N$. It remains to show that $h|_{S^{v}}$ is homotopic to $g$ as a pointed map. Here we follow the last paragraph of the proof of [BDPW23, Theorem 5.2.3]. Namely, given $u\in S^{v}$, suppose $u$ is in $U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$ and not in $U_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{0},\ldots,\lambda_{n}\}$. Associated to $g(u)$ is a sequence of $\kappa$ elements in $\operatorname{hur}^{X}$, for some integer $\kappa$. We can first linearly stretch these $\kappa$ elements each by a factor of $\frac{1}{2r}$, and then use the identifications of (3) of the 2-sided bar construction § A.3 to make $n$ cuts in the places specified by $t_{\lambda_{i}}$ in $\sigma\circ g$. Let $\tilde{g}$ be the function of $u$ obtained this way, so that $\tilde{g}$ is homotopic to $g$. Then, $\tilde{g}$ agrees with $h|_{S^{v}}$ except for the weights defining a point in $\Delta^{n}$. Linearly changing these weights then gives a continuous homotopy between $\tilde{g}$ and $h|_{S^{v}}$. Thus $g$ is homotopic to $h|_{S^{v}}$. ∎
@@ A.4. A simple topological model
Let $M$ be a set with a right action of $\operatorname{Hur}^{X}$ and let $N$ be a set with a left action of $\operatorname{Hur}^{X}$. Let $10<\epsilon<$ 1. We define $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ to be the set consisting of triples $(m,n,x,l:x\to X)$ with $m\in M,n\in N,$ $x=\{p_{1},\ldots,p_{k}\}$ a finite subset of $[0,1]\times[\epsilon,1-\epsilon]$ (with first projection $\pi_{1}:[0,1]\times[0,1]\to[0,1]$ and second projection $\pi_{2}:[0,1]\times[0,1]\to[0,1]$) such that for each $i\neq j$, the distance between $\pi_{2}(p_{i})$ and $\pi_{2}(p_{j})$ is at least $\epsilon$, and $l:x\to X$ is a map of sets.
@@
    Suppose $x=\{p_{1},\ldots,p_{v},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{v})=0$, so $p_{r}$ lies on the left boundary, the data ($m,n,x,l:x\to X)$ is equivalent to $(m^{\prime},n,x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to X)$, defined as follows: write $(y_{1},\ldots,y_{r}):=\sigma_{1}\sigma_{2}\cdots\sigma_{r-1}(l(p_{1}),\ldots,l(p_{r})).$ Then $m^{\prime}=my_{1}$, and $l^{\prime}(p_{i}):=y_{i}$ for $1\leq i<r$, $l^{\prime}(p_{i}):=l(p_{i})$ for $i>r$.
@@
For technical reasons, we will want to work with inductive systems of topological spaces $(X_{\epsilon})_{\epsilon>0}$, which we view as presenting the colimit of their homotopy types. For example, we will consider $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ as an ind-system as $\epsilon$ approaches 00.
@@
[…]

### v3

@@ A.3. A scanning map
Next, we produce well behaved topological models for the source and target of § 4.5. The following definition will be the key ingredient in constructing the models. The reader may wish to consult § A.3 for help understanding § A.3.
@@
Let $M$ be a set with a right action of $\operatorname{Hur}^{X}$ and $N$ be a set with a left action of $\operatorname{Hur}^{X}$. Consider the topological space $B[M,\operatorname{Hur}^{X},N]$ consisting of points which are of the form
@@
    A finite collection of pairwise disjoint open balls $U_{1},\ldots,U_{n}$ in contained in the interior of $d_{1},d_{2}]\times[0,1]$.
@@
    A homotopy class of paths $\phi$ from the configuration of the centers of the balls $U_{i}$, viewed as an element of $\mathrm{confbig}_{n}$, to the base point.
@@
    None of the points in $x$ lie in $d_{1},d_{2}]\times[0,1]-\cup_{1}^{n}U_{i}$, and there is a unique point from $x$ in each $U_{i}$.
@@
    Cutting the element of $\operatorname{hurbig}^{X}$ twice to restrict it to the interval $d_{1},d_{2}]$ yields a point $y^{\prime}\in\operatorname{hurbig}^{X}$ (see § A.2). Then, using the homotopy class of $\phi$, the corresponding tuple of elements in $X$ associated to $y^{\prime}$ is $\alpha_{1},\ldots,\alpha_{n}$.
@@
    Define $y_{1}\in\operatorname{hurbig}^{X}$ to be the element of $\operatorname{hur}^{X}$ obtained by cutting and restricting to the interval $0,d_{1}]$, and let $y_{2}$ be the element similarly obtained by using the interval $d_{2},1]$. We then require that $ay_{1}=a^{\prime}$ and $y_{2}b=b^{\prime}$.
@@
The topology of § A.3 is such that points of the configuration $x$ are allowed to collide with the boundary $\{0,1\}\times[0,1]$, at which point they disappear and multiply the elements $m,n$ using the left action of $\operatorname{Hur}^{X}$ on $M$ and the right action on $N$. Condition (3) guarantees that when a collection of points converges to the boundary, it multiplies $m$ or $n$ to the correct value.
We next construct a homotopy equivalence showing that the topological space $B[M,\operatorname{Hur}^{X},N]$ models the two-sided bar construction $M\otimes_{\operatorname{hur}^{X}}N$, which we define next. Our proof is an adaptation of the proof of [BDPW23, Theorem 5.2.3]. Indeed, the case M=N=X=M=N=X=* corresponds to the result there.
@@
We produce a map $\sigma:M\otimes_{\operatorname{hur}^{X}}N\to B[M,\operatorname{Hur}^{X},N]$, as follows: Given data as in (A.2) for $H=\operatorname{hur}^{X}$, and using notation from § A.2, let $t_{i}:=t(x_{i})\in\mathbb{R}_{\geq 0}$ denote the time associated to $x_{i}\in\operatorname{hur}^{X}$. We multiply $x_{1},\ldots,x_{n}$ to give an element $x\in\operatorname{hur}^{X}$ with $t:=t(x)=\sum_{i=1}^{n}t_{i}$. First, extend the interval on which $x$ is defined from $0,t]\times[0,1]$ to $\frac{1}{2},t+\frac{1}{2}]\times[0,1]$. We then define $t^{\prime}:=\sum_{i=1}^{n}y_{i}\cdot(\sum_{j=1}^{i}t_{j})$. At this point, it will be useful to recall the cutting construction from § A.2. Choose $\epsilon>0$ small enough so that there are no points of the configuration $x$ in $(t^{\prime}-\frac{1}{2},t^{\prime}-\frac{1}{2}+\epsilon]\times[0,1]$ and $t^{\prime}+\frac{1}{2}-\epsilon,t^{\prime}+\frac{1}{2})\times[0,1]$. We can then cut $x$ into three pieces, $w,x^{\prime\prime},z\in\operatorname{hurbig}^{X}$ coming from the intervals $\frac{1}{2},t^{\prime}-\frac{1}{2}+\epsilon]$, $t^{\prime}-\frac{1}{2}+\epsilon,t^{\prime}+\frac{1}{2}-\epsilon],[t^{\prime}+\frac{1}{2}-\epsilon,t+\frac{1}{2}]$. Extend $x^{\prime\prime}$ to be of length $1$ by extending the length of the interval on both ends by $\epsilon$, and let $x^{\prime}$ denote the resulting element of $\operatorname{hurbig}^{X}$. Note that $x^{\prime}$ doesn’t depend on $\epsilon$, and neither does the class of $w,z$ in $\pi_{0}\operatorname{Hur}^{X}$.
@@
Having shown $\sigma$ is continuous, we next claim that $\sigma$ is surjective on path components. Every point in $B[M,\operatorname{Hur}^{X},N]$ has a path to a point where the configuration on $0,1]\times[0,1]$ is empty, by linearly pushing the configuration towards $0\times[0,1]$ on the boundary, and multiplying once points collide with the boundary. A point with an empty configuration is determined by $(a,b)\in M\times N$. This point is the image of the point $(a,b,(1),())\in M\otimes_{\operatorname{hur}^{X}}N$, showing the claimed surjectivity on path components. We view these points as above with empty configuration as giving a copy of $M\times N\subset M\otimes_{\operatorname{hur}^{X}}N$.
@@
We turn to proving $(i)$, and assume that we have a map $f$ as given in the statement of $(i)$. We set up notation to define the map $g$ described in $(i)$. Using compactness, we can find a finite open cover $D^{v}=\cup_{\lambda\in\Lambda}U_{\lambda}$ such that there is an $r>0$ and for each $\lambda\in\Lambda$, there is a $t_{\lambda}\in(0,1)$ such that for all $u\in U_{\lambda}$, the configuration associated to $f(u)$ has no points in $(t_{\lambda}-r,t_{\lambda}+r)\times[0,1]$. Choose a partition of unity $w_{\lambda}$ subordinate to the cover $U_{\lambda}$. Let $u$ be a point in $U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$, and not in $U_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{0},\ldots,\lambda_{n}\}$. Note that for every point $u\in D^{v}$ satisfies the above for a unique choice of $\{\lambda_{0},\ldots,\lambda_{n}\}\subset\Lambda$ by taking $\lambda_{0},\ldots,\lambda_{n}$ to be the subset of $\Lambda$ such that $u\in U_{\lambda_{i}}$ for every such $i$. Assume that $t_{\lambda_{0}}<t_{\lambda_{1}}<\cdots<t_{\lambda_{n}}$ (which can always be achieved by decreasing $r$ and making all $t_{\lambda}$ distinct). Then we consider the $n$-tuple of elements $x_{1},\ldots,x_{n}\in\operatorname{hur}^{X}$, where $x_{i}$ is obtained from the element of $\operatorname{hur}$ associated to $f(u)$ by cutting along $t_{\lambda_{i-1}}\times[0,1]$ and $t_{\lambda_{i}}\times[0,1]$, taking the configuration with points lying in $t_{\lambda_{i-1}},t_{\lambda_{i}}]\times[0,1]$, and rescaling the intervals by $\frac{1}{2r}$ so that the time function $t$ applied to this new configuration is $(t_{\lambda_{i}}-t_{\lambda_{i-1}})\frac{1}{2r}$. Define $u^{\prime}$ and $u^{\prime\prime}\in\operatorname{hurbig}^{X}$ to be the elements obtained respectively by cutting to restrict to the intervals $0,t_{\lambda_{0}}]$ and $t_{\lambda_{n}},1]$ respectively. We define the map $g:D^{v}\to M\otimes_{\operatorname{hur}^{X}}N$ by sending $u$ to the tuple
@@
Choose a basic open set $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$ containing $f(u)$ and a small open neighborhood $V$ containing $u$ with $V\subset f^{-1}(\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime}))\cap U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$. We may shrink $\mathfrak{B}(d_{1},d_{2},U_{i},\phi,\alpha_{i},a^{\prime},b^{\prime})$, to assume $d_{1}\leq t_{\lambda_{0}},d_{2}\geq t_{\lambda_{n}}$, where, here, shrinking this neighborhood may involve increasing the number of $U_{\lambda}$ so that there is one such open containing each point of the configuration $x$, appearing in the data for $f(u)$, in $(d_{1},d_{2})\times[0,1]$. By further shrinking $V$, we may assume the functions $w_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{0},\ldots,\lambda_{n}\}$ vanish on $V$. Therefore, using identifications (1)(3)(1)-(3) of the bar construction, for the purposes of proving continuity, we may assume our open cover consists solely of $U_{\lambda_{0}},\ldots,U_{\lambda_{n}}$, i.e, we can ignore the other $U_{\lambda}$. At this point, condition (3) of § A.3 enables us to further shrink $V$ so that ($\star$) holds for all $g(v)$ with $v\in V$. We have therefore found an open set in $D^{v}$ containing $u$ and contained in $g^{-1}(W)$ for $W$ an open set containing $g(u)$. This proves that $g$ is continuous.
@@
To prove $(ii)$, given a map of pairs $g:(S^{v},*)\to(M\otimes_{\operatorname{hur}^{X}}N,M\times N)$ and a nullhomotopy of $\sigma\circ g$, the same construction used above for $(i)$ lifts the nullhomotopy of $\sigma\circ g$ to a pointed map $h:D^{v+1}\to M\otimes_{\operatorname{hur}^{X}}N$. It remains to show that $h|_{S^{v}}$ is homotopic to $g$ as a pointed map. Here we follow the last paragraph of the proof of [BDPW23, Theorem 5.2.3]. Namely, given $u\in S^{v}$, suppose $u$ is in $U_{\lambda_{0}}\cap\cdots\cap U_{\lambda_{n}}$ and not in $U_{\lambda}$ for $\lambda\in\Lambda-\{\lambda_{0},\ldots,\lambda_{n}\}$. Associated to $g(u)$ is a sequence of $\kappa$ elements in $\operatorname{hur}^{X}$, for some integer $\kappa$. We can first linearly stretch these $\kappa$ elements each by a factor of $\frac{1}{2r}$, and then use the identification (3) of the 2-sided bar construction § A.3 to make $n$ cuts in the places specified by $t_{\lambda_{i}}$ in $\sigma\circ g$. Let $\tilde{g}$ be the function of $u$ obtained this way, so that $\tilde{g}$ is homotopic to $g$. Then, $\tilde{g}$ agrees with $h|_{S^{v}}$ except for the weights defining a point in $\Delta^{n}$. Linearly changing these weights then gives a continuous homotopy between $\tilde{g}$ and $h|_{S^{v}}$. Thus $g$ is homotopic to $h|_{S^{v}}$. ∎
@@ A.4. A simple topological model
Let $M$ be a set with a right action of $\operatorname{Hur}^{X}$ and let $N$ be a set with a left action of $\operatorname{Hur}^{X}$. Let $0<\epsilon<1$. We define $Q_{\epsilon}[M,\operatorname{Hur}^{X},N]$ to be the set consisting of triples $(m,n,x,l:x\to X)$ with $m\in M,n\in N,$ $x=\{p_{1},\ldots,p_{k}\}$ a finite subset of $0,1]\times[\epsilon,1-\epsilon]$ (with first projection $\pi_{1}:[0,1]\times[0,1]\to[0,1]$ and second projection $\pi_{2}:[0,1]\times[0,1]\to[0,1]$) such that for each $i\neq j$, the distance between $\pi_{2}(p_{i})$ and $\pi_{2}(p_{j})$ is at least $\epsilon$, and $l:x\to X$ is a map of sets.
@@
    Suppose $x=\{p_{1},\ldots,p_{r},\ldots,p_{k}\}$ with $\pi_{2}(p_{1})>\cdots>\pi_{2}(p_{k})$. If $\pi_{1}(p_{r})=0$, so $p_{r}$ lies on the left boundary, the data ($m,n,x,l:x\to X)$ is equivalent to $(m^{\prime},n,x^{\prime}:=\{p_{1},\ldots,p_{r-1},p_{r+1},\ldots,p_{k}\},l^{\prime}:x^{\prime}\to X)$, defined as follows: write $(y_{1},\ldots,y_{r}):=\sigma_{1}\sigma_{2}\cdots\sigma_{r-1}(l(p_{1}),\ldots,l(p_{r})).$ Then $m^{\prime}=my_{1}$, and $l^{\prime}(p_{i}):=y_{i}$ for $1\leq i<r$, $l^{\prime}(p_{i}):=l(p_{i})$ for $i>r$.
@@
Construction A.4.0.
Choose a smooth function $f:[0,1]\to\mathbb{R}$ such that $f([0,\frac{1}{2}))<0$ and f(1x)=f(x)f(1-x)=-f(x). In particular, $f(1/2)=0$. Consider the vector field $f(x)\frac{\partial}{\partial x}$ on $0,1]\times[0,1]$, where $x$ is the first coordinate. We can then form a continuous flow $\phi_{t}:B[M,\operatorname{Hur}^{X},N]\times[0,\infty)\to B[M,\operatorname{Hur}^{X},N]$ by moving points in a configuration along the flow, and having them act on the elements $(m,n)$ in $M\times N$ when they reach the boundary. We can define such a flow (which we give the same name) on $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ via the same procedure.
Proposition A.4.0.
Figure 2. This is a picture depicting the compatibility of the map $\tilde{g}$ in the proof of § A.4 with relations $(i)$ and $(ii)$ from § A.4. The first two columns depict the paths $\gamma$ and $\gamma^{\prime}$ the points take during the map $\tilde{g}$ under relation $(i)$, and the composite $\gamma^{\prime}\gamma^{-1}$ is depicted underneath. The last two columns depict the paths $\gamma$ and $\gamma^{\prime}$ which the points take under relation $(ii)$, and the composite $\gamma^{\prime}\gamma^{-1}$ is again depicted underneath. In the diagrams, there are 44 points, $p_{1}$ which is purple, $p_{2}$ which is red, $p_{3}$ which is green, and $p_{4}$ which is blue. So $r=2$ and $k=4$ in the notation of § A.4. The point $p_{2}$ is on the boundary. When $p_{2}$ is on the left boundary, the path is $\sigma_{1}\cdots\sigma_{r-1}=\sigma_{1}$ since $r=2$. When $p_{2}$ is on the right boundary, the path is $\sigma_{r}\cdots\sigma_{k-1}=\sigma_{2}\sigma_{3}$.
Proof.
$Q\to B[M,\operatorname{Hur}^{X},N]$
We define the interior $Q^{\circ}$ to be the dense subset of $Q$ on which the configurations are contained in $(0,1)\times[\epsilon,1-\epsilon]$. In other words, a point of $Q^{\circ}$ is given by $m\in M,n\in N$, such a configuration $x\subset(0,1)\times[\epsilon,1-\epsilon]$, and a function $l:x\to X$. We define a continuous function $g:Q^{\circ}\to B[M,\operatorname{Hur}^{X},N]$ as we describe next.
We first continuously choose a homotopy class of paths $\gamma$ from the configuration $x$ to $\mathrm{bigord}$ as follows: given a configuration $x$, because the projection $\pi_{2}|_{x}:x\to[0,1]$ to the second coordinate is injective, we can linearly move all points so that their first coordinate is $\frac{1}{2}$. Then, we can rotate the configuration in $0,1]\times[0,1]$ counterclockwise by $\frac{\pi}{2}$ so that the second coordinate of all the points is $\frac{1}{2}$, so that the path ends in $\mathrm{bigord}$. This ends our description of $\gamma$.
We claim the above map $Q^{\circ}\to B[M,\operatorname{Hur}^{X},N]$ continuously extends to a map $Q\to B[M,\operatorname{Hur}^{X},N]$. To see this, we can produce a continuous inward flow $\varphi_{-t}$ on $Q$ by pushing points inwards along the vector field $f(x)\frac{\partial}{\partial x}$ as in § A.4. For positive $t$, this flow lands inside $Q^{\circ}$, so we can continuously extend the function $g$ to a map $\tilde{g}:Q\to B[M,\operatorname{Hur}^{X},N]$ via the formula $\phi_{t}\circ g\circ\varphi_{-t}$ for any $t>0$, with $\phi_{t}$ as in § A.4. It is easy to see that the underlying configuration associated to the map $\tilde{g}$ is similar to the configuration of the original point, with the only difference being that all of the points on the boundary of $0,1]\times[0,1]$ are removed.
We now establish relation $(ii)$ from § A.4 in this case, as $(i)$ follows from a similar argument. Suppose that we have a point $w\in Q$ where $y$ is the only point in the configuration associated to $w$ on the boundary of $0,1]\times[0,1]$. Moreover, assume $y$ lies on $1\times[0,1]$. We need to compute the limit of $\phi_{s}\circ g\circ\varphi_{-t}(w)$ as $s$ approaches $t$ from below. For the configuration associated to point $\varphi_{-t}(w)$, we can choose some $\delta<1$ so that the point $y^{\prime}$ to which $y$ flows in in $\varphi_{-t}(w)$ has $\pi_{1}(y^{\prime})>\delta$ (where $\pi_{1}(y^{\prime})$ denotes the first coordinate of $y^{\prime}$) and is the only point of the configuration with first coordinate $\geq\delta$. Let $\gamma]$ be the homotopy class of paths from the configuration in $\varphi_{-t}(w)$ to $\mathrm{bigord}$ that is used in defining the map $g$. We can choose a different homotopy class of paths by first cutting up $0,1]\times[0,1]$ into $0,\delta]\times[0,1]$ and $\delta,1]\times[0,1]$, doing the construction as in the map $g$ on each piece [^6] Really, the rotation used in the construction should be rescaled since $0,\delta]\times[0,1]$ and $\delta,1]\times[0,1]$ are not squares., and concatenating to get a new homotopy class of paths $\gamma^{\prime}]$. One checks that the class $\gamma^{\prime}][\gamma]^{-1}$ in the braid group is given by $\sigma_{k-1}\cdots\sigma_{i}$, where there are $k$ points of the configuration, and where $y$ has the $i$th highest second coordinate of all elements in the configuration. We have illustrated checking this in Figure 2. Thus $g\circ\varphi_{-t}(w)$ is equivalent to the point where the path $\gamma^{\prime}]$ is used instead of $\gamma]$, and the element of $X^{k}$ is changed according to the action of $\sigma_{k-1}\cdots\sigma_{i}$. Letting $s$ approach $t$, we see that this is exactly the image of the point as described in relation $(ii)$.
@@
For technical reasons, we will want to work with inductive systems of topological spaces $(X_{\epsilon})_{\epsilon>0}$, which we view as presenting the colimit of their homotopy types. For example, we will consider $\overline{Q}_{\epsilon}[M,\operatorname{Hur}^{X},N]$ as an ind-system as $\epsilon$ approaches 0.
@@
    The restriction of $H$ along $i\wedge\operatorname{id}$ induces a pointed homotopy $K^{\prime}\wedge[0,1]_{+}\to X^{\prime}_{\beta}$ whose restriction to the point $0\in[0,1]_{+}$ is the composite $K^{\prime}\xrightarrow{\psi}X^{\prime}_{\alpha}\to X^{\prime}_{\beta}$ where $\psi$ is the map from (A.3).